OpenStudy (anonymous):

find the linear function f(x) for which both f(-2) and f(0)=-4

jimthompson5910 (jim_thompson5910):

If you have a linear function that has 2 x values produce the same y value, then this function must graph a horizontal line.

OpenStudy (anonymous):

it's f(-2)=6, sorry

jimthompson5910 (jim_thompson5910):

oh nvm

jimthompson5910 (jim_thompson5910):

are you able to find the slope through the two points (-2,6) and (0,-4) ?

OpenStudy (anonymous):

How do I come up with that?

jimthompson5910 (jim_thompson5910):

are you familiar with the slope formula?

OpenStudy (anonymous):

It's y=mx+b right?

jimthompson5910 (jim_thompson5910):

that's slope intercept form, which is close to what I'm referring to

jimthompson5910 (jim_thompson5910):

I'm actually talking about this formula \[\Large m = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

I don't know which one to plug which into. I think that's my problem.

jimthompson5910 (jim_thompson5910):

Basically you find the y difference by subtracting the y coordinates Do the same for the x coordinates (in the same order) then you divide the y difference by the x difference

jimthompson5910 (jim_thompson5910):

(x1,y1) is your first point (x2,y2) is the second

jimthompson5910 (jim_thompson5910):

In this case (x1,y1) = (-2,6) (x2,y2) = (0,-4)

OpenStudy (anonymous):

So it would look like -4-6 over 0-2?

jimthompson5910 (jim_thompson5910):

very close, but it should be -4-6 over 0 - (-2)

OpenStudy (anonymous):

So the slope would be -5/1 ?

jimthompson5910 (jim_thompson5910):

which reduces to -5

jimthompson5910 (jim_thompson5910):

the slope is m = -5 the y-intercept is b = -4 ... this is because f(0) = -4

jimthompson5910 (jim_thompson5910):

plug m = -5 and b = -4 into y = mx + b to get y = mx + b y = -5x + (-4) y = -5x - 4

jimthompson5910 (jim_thompson5910):

Therefore, the linear function is f(x) = -5x - 4

OpenStudy (anonymous):

Awesome. I think I get it now.

jimthompson5910 (jim_thompson5910):

That's great that you do